Consider this system at equilibrium. A(aq) forward and reverse arrows B (aq)      Delta H= -650 kj/mol

What can be said about Q and K immediately after an increase in temperature?
Answer #1

given\; A_{\left ( aq \right )}\Leftrightarrow B_{\left ( aq \right )}\; \Delta H=-650kj/mol

since \Delta H=-ve increase in temperature moves the equilibrium in backward direction

so Q = \frac{\left [ A \right ]}{\left [ B \right ]} \; and \; K=K^{-1}